# Principal twistor models and asymptotic hyperkähler metrics

### 04 Mar 2026 v1

@@ 4.1.2. Injectivity theorem for the moduli space
By Lemma 4.7, under multiplication by $\varepsilon>0$ for the fiberwise $\mathbb{C}^{*}$-action $\tilde{\lambda}$ on the principal twistor model, the family $\{\tilde{\lambda_{\varepsilon}}\circ\ell_{y,s}\}_{y\in Y}$ forms a family of twistor lines on the twistor model $Z_{\varepsilon^{2}s}$. Thus, we define $\ell_{y,\varepsilon^{2}s}$ as the member of this family passing through the point $y\in Y$, and consider the family of twistor lines $\{\ell_{y,\varepsilon^{2}s}\in\mathcal{T}_{\varepsilon^{2}s}\}_{y\in Y,\varepsilon>0}$. For this family, the assertion of the lemma holds. ∎
@@
Suppose that $X$ is a conical symplectic variety with a crepant resolution $Y$, and consider the universal Poisson deformation $\mathcal{Y}\to\mathcal{C}$ of $Y$. Let $f:\mathcal{Y}\to Y\times\mathcal{C}$ be the differentiable trivialization obtained in Lemma 4.3. Denoting by $E$ the exceptional set of $Y$ and setting $\mathcal{E}:=f^{-1}(E\times\mathcal{C})$, the open submanifold $\mathcal{Y}\setminus\mathcal{E}$ admits an $\mathbb{R}^{+}$-equivariant differentiable trivialization $f^{\prime}:\mathcal{Y}\setminus\mathcal{E}\to(Y\setminus E)\times\mathcal{C}$. Here, the target space $(Y\setminus E)\times\mathcal{C}$ is equipped with the $\mathbb{C}^{*}$-action $\lambda_{Y}$ on $Y$ and the scalar multiplication of weight $2$ on $\mathcal{C}$.
@@
Here, we used the relation $\tilde{\lambda}_{Y,r}^{*}d_{0}=r^{2}d_{0}$. Therefore, considering the family of twistor lines $\{\ell_{y,\varepsilon^{2}s}\in\mathcal{T}_{\varepsilon^{2}s}\}_{y\in(Y\setminus E),\varepsilon\geq 0}$ extended to $\varepsilon\geq 0$, it follows that this family gives a lift of the set $(Y\setminus E)\times\{\varepsilon^{2}s\mid\varepsilon\geq 0\}$ with respect to the covering map $\Phi$ in Lemma 4.5. That is, the family of twistor lines corresponding to the metric $g_{s}$ and the family of twistor lines corresponding to the metric $g_{0}$ exist on the same sheet with respect to the covering map $\Phi$. From this, the assertion of the proposition follows. ∎
@@
Let $X$ be a conical symplectic variety such that $X_{\mathrm{reg}}$ admits an algebraic hyperkähler cone metric $g_{0}$, and let $Y$ be a crepant resolution of $X$. Let $\mathcal{Y}(1)\to\mathcal{C}(2)$ be the principal twistor model constructed in Theorem 3.39. Consider the moduli space $\mathcal{M}$ of hyperkähler structures defined by
@@ 4.1.3. Openness of Twistor Space in the Principal Twistor Model
Let $X$ be a conical symplectic variety. Assume that $X_{\mathrm{reg}}$ admits an algebraic hyperkähler cone metric $g_{0}$. Let $Y$ be a crepant resolution of $X$. In this subsection, we show that a certain kind of openness of the twistor space in the principal twistor model holds when $X$ has an isolated singularity.
@@
Let $X$ be a conical symplectic variety with an isolated singularity, and assume that $X_{\mathrm{reg}}$ admits an algebraic hyperkähler cone metric $g_{0}$. Let $Y$ be a crepant resolution of $X$. Let $(\mathcal{Y}(1)\to\mathcal{C}(2),\omega_{P},\tau_{P})$ be the principal twistor model given by Theorem 3.39, and let $H^{0}(\mathbb{P}^{1},\mathcal{C}(2))^{\sigma_{2}}$ denote the real vector space of real sections of the vector bundle $\mathcal{C}(2)$. For a real section $s\in H^{0}(\mathbb{P}^{1},\mathcal{C}(2))^{\sigma_{2}}$, let $Z_{s}$ be the twistor model obtained by slicing $\mathcal{Y}(1)$ along $s$. If $Z_{s}$ is the twistor space associated with a hyperkähler metric on $Y$ that is asymptotic to $g_{0}$ at infinity, then for any real section $s^{\prime}$ in a sufficiently small open neighborhood of $s$, the model $Z_{s^{\prime}}$ is also the twistor space of such a metric.
@@
    The twistor model $Z_{s^{\prime}}$ is a twistor space: Consider a relatively compact open set $U$ containing both the real section $s\in H^{0}(\mathbb{P}^{1},\mathcal{C}(2))^{\sigma_{2}}$ and the zero section. By Assertion (1), there exists a compact set $K$ containing the exceptional set $E$ on the resolution $Y$ such that the map $\Phi$ is surjective onto the set $(Y\setminus K)\times U$. On the other hand, since $Z_{s}$ is a twistor space by assumption, for the compact set $K$ on $Y$, by choosing a sufficiently small open neighborhood $U^{\prime}\subset U$ of the real section $s\in H^{0}(\mathbb{P}^{1},\mathcal{C}(2))^{\sigma_{2}}$, the map $\Phi$ is surjective onto the set $K\times U^{\prime}$. Therefore, the map $\Phi$ is surjective onto the set $Y\times U^{\prime}$. Now, we may assume that the neighborhood $U^{\prime}$ is simply connected, and the crepant resolution $Y$ is also simply connected. This allows us to lift the set $Y\times U^{\prime}$ to the finite covering $\Phi$, and we can choose the lift that contains the twistor lines of $Z_{s}$. Thus, for any real section $s^{\prime}\in U^{\prime}$, the twistor model $Z_{s^{\prime}}$ is a twistor space.
@@
As a corollary of this proposition, we obtain the following result on the moduli of hyperkähler structures with asymptotic behavior.
@@
Let $X$ be a conical symplectic variety with an isolated singularity. Assume that $X_{\mathrm{reg}}$ admits an algebraic hyperkähler cone metric $g_{0}$. Let $Y$ be a crepant resolution of $X$. Then, the following two assertions hold:
@@
In particular, if the moduli space $\mathcal{M}$ (resp. $\mathcal{M}_{\text{met}}$) is non-empty, then we have $\dim_{\mathbb{R}}\mathcal{M}=3d$ (resp. $\dim_{\mathbb{R}}\mathcal{M}_{\text{met}}=3d-3$). Here, $d=\dim H^{2}(Y;\mathbb{C})$.
@@
Corollary 4.18 applies to the closures of minimal nilpotent orbits in complex semisimple Lie algebras. Let $X=\overline{\mathcal{O}_{\mathrm{min}}}$ be the closure of the minimal nilpotent orbit (cf. [3]) in a complex semisimple Lie algebra $\mathfrak{g}$. The space $X$ is a conical symplectic variety with an isolated singularity at the origin, and its regular locus $X_{\mathrm{reg}}=\mathcal{O}_{\mathrm{min}}$ admits an exact hyperkähler cone metric via Nahm’s equations, as constructed by Kronheimer [10]. As a specific example, consider the case $\mathfrak{g}=\mathfrak{sl}_{n}(\mathbb{C})$ for $n\geq 2$, where $X$ admits a crepant resolution $Y=T^{*}\mathbb{P}^{n-1}$. On this resolution, there exists the Calabi metric, which is an asymptotically conical hyperkähler metric that approaches the cone metric at infinity. In this setting, Corollary 4.18 shows that the dimension of the moduli space $\mathcal{M}_{\text{met}}$ of such asymptotic metrics is exactly $3\dim H^{2}(T^{*}\mathbb{P}^{n-1};\mathbb{C})-3=0$ (meaning the metric is isolated up to scaling).
@@ 4.2.1. Metrics Constructed by Hyperkähler Quotients
    $d\langle\mu_{A},\xi\rangle(v)=\omega_{A}(v,\xi^{*}).$
    Here, $\langle\cdot,\cdot\rangle$ denotes the pairing between $\mathfrak{g}$ and $\mathfrak{g}^{*}$, and $\omega_{A}\coloneqq g(A\cdot,\cdot)$.
@@
By applying the theory of symplectic quotients and geometric invariant theory by Kempf-Ness and Kirwan, the following lemma is obtained (cf. [11, Theorem 3.1, Theorem 4.1]). The proof is omitted.
@@
    If $X_{\zeta}$ is non-singular, it naturally inherits a hyperkähler structure. Since the map $\pi_{I}$ preserves the descended $I$-holomorphic symplectic form, it automatically gives a crepant resolution of $X_{\zeta_{0}}$.
@@
    The hyperkähler quotient $X_{\zeta}$ is naturally diffeomorphic to a crepant resolution $Y$ of $X_{0}$: We follow the argument of Kronheimer [10]. Let $g_{\zeta}$ be the hyperkähler metric on the hyperkähler quotient $X_{\zeta}$. Here, we denote the point $\zeta$ by $\zeta=(\zeta_{I},\zeta_{J},\zeta_{K})\in\mathbb{R}^{3}\times\mathfrak{g}^{*}$. Since the group $G$ acts linearly on the vector space $M$, the set of points $\xi\in\mathbb{R}^{3}\times\mathfrak{g}^{*}$ such that $X_{\xi}\coloneqq\mu^{-1}(\xi)/G$ admits singularities can be written as a union of finitely many hyperplanes in the vector space $\mathbb{R}^{3}\times\mathfrak{g}^{*}$. (That is, it has a wall-chamber structure.) Therefore, by considering the $SO(3)$-action on $\mathbb{R}^{3}$ by hyperkähler rotation, we may assume that the parameter $\zeta$ is generic so that $X_{\zeta}$, $X_{\zeta^{\prime}}$, and $X_{\zeta^{\prime\prime}}$ are all non-singular, where we set $\zeta^{\prime}\coloneqq(0,\zeta_{J},\zeta_{K})$ and $\zeta^{\prime\prime}\coloneqq(0,0,\zeta_{J})$. Because these spaces are non-singular, the maps $\pi_{I},\pi_{J}$, and $\pi_{K}$ obtained in Lemma 4.24 automatically give crepant resolutions (see Remark 4.25), and we obtain a smooth map
@@
    as their composition. Now, since these intermediate spaces are non-singular by our generic choice, in the map (4.3), $\pi_{I}$ and $\pi_{J}$ are diffeomorphisms. Therefore, if we let $Y\coloneqq X_{\zeta^{\prime\prime}}$ be the $K$-holomorphic crepant resolution of $X_{0}$, $X_{\zeta}$ is naturally diffeomorphic to the crepant resolution $Y$.
@@
By the above proposition, Theorem 3.39 can be applied to Nakajima quiver varieties (cf. Nakajima [11]) and toric hyperkähler manifolds (cf. Bielawski and Dancer [1]). Note that for toric hyperkähler manifolds, this application requires the hyperkähler moment map to satisfy $\mu(0)=0$. Under this condition—which is automatically satisfied for quiver varieties by definition—one can directly verify from the explicit forms of the moment maps that the required condition $\lambda_{c}^{*}\mu=|c|^{2}\mu$ holds in both cases.
@@ 4.2.2. QALE Hyperkähler Metrics
Let $G<Sp(n)$ be a finite group, and let $X=\mathbb{C}^{2n}/G$ be the quotient space. Assume that $X$ admits a crepant resolution $Y$. We assume that any QALE hyperkähler metric on $Y$ (cf. [7, Definition 2.1]), provided it exists, is necessarily algebraic.
@@
According to Joyce [7, p.115], the above is proved by hypercomplex algebraic geometry, but the details are not described. Note that it is not known whether a QALE hyperkähler metric exists on an arbitrary crepant resolution $Y$.
@@
$X$ has an isolated singularity if and only if $\dim_{\mathbb{C}}Y=2$. In this case, a QALE hyperkähler metric on $Y$ coincides with an ALE gravitational instanton. Thus, by Corollary 4.18, we deduce that the moduli space of ALE gravitational instantons has real dimension 3d33d-3. Here, $d=\dim H^{2}(Y;\mathbb{C})$. This result also follows immediately from the result of Kronheimer [9].

### 12 Mar 2026 v2

@@ 4.1.2. Injectivity theorem for the moduli space
By Lemma 4.7, under multiplication by $\varepsilon>0$ for the fiberwise $\mathbb{C}^{*}$-action $\tilde{\lambda}$ on the principal twistor model, we obtain the family of twistor lines $\{\tilde{\lambda_{\varepsilon}}\circ\ell_{y,s}\}_{y\in Y}$ on the twistor model $Z_{\varepsilon^{2}s}$. Thus, we define $\ell_{y,\varepsilon^{2}s}$ as the member of this family passing through the point $y\in Y$, and consider the family of twistor lines $\{\ell_{y,\varepsilon^{2}s}\in\mathcal{T}_{\varepsilon^{2}s}\}_{y\in Y,\varepsilon>0}$. For this family, the assertion of the lemma holds. ∎
@@
Suppose that $X$ is a conical symplectic variety with a crepant resolution $Y$, and consider the universal Poisson deformation $\mathcal{Y}\to\mathcal{C}$ of $Y$. Let $f:\mathcal{Y}\to Y\times\mathcal{C}$ be the differentiable trivialization obtained by Lemma 4.3. Denoting by $E$ the exceptional set of $Y$ and setting $\mathcal{E}:=f^{-1}(E\times\mathcal{C})$, the open submanifold $\mathcal{Y}\setminus\mathcal{E}$ admits an $\mathbb{R}^{+}$-equivariant differentiable trivialization $f^{\prime}:\mathcal{Y}\setminus\mathcal{E}\to(Y\setminus E)\times\mathcal{C}$. Here, the target space $(Y\setminus E)\times\mathcal{C}$ is equipped with the $\mathbb{C}^{*}$-action $\lambda_{Y}$ on $Y$ and the scalar multiplication of weight $2$ on $\mathcal{C}$.
@@
Here, we used the relation $\tilde{\lambda}_{Y,r}^{*}d_{0}=r^{2}d_{0}$. Therefore, considering the family of twistor lines $\{\ell_{y,\varepsilon^{2}s}\in\mathcal{T}_{\varepsilon^{2}s}\}_{y\in(Y\setminus E),\varepsilon\geq 0}$ extended to $\varepsilon\geq 0$, it follows that this family gives a lift of the set $(Y\setminus E)\times\{\varepsilon^{2}s\mid\varepsilon\geq 0\}$ with respect to the covering map $\Phi$ in Lemma 4.5. That is, the families of twistor lines corresponding to the metrics $g_{s}$ and $g_{0}$ exist on the same sheet with respect to the covering map $\Phi$. From this, the assertion of the proposition follows. ∎
@@
Let $X$ be a conical symplectic variety such that the regular locus $X_{\mathrm{reg}}$ admits an algebraic hyperkähler cone metric $g_{0}$, and let $Y$ be a crepant resolution of $X$. Let $\mathcal{Y}(1)\to\mathcal{C}(2)$ be the principal twistor model constructed in Theorem 3.39. Consider the moduli space $\mathcal{M}$ of hyperkähler structures defined by
@@ 4.1.3. Openness of Twistor Space in the Principal Twistor Model
Let $X$ be a conical symplectic variety. Assume that the regular locus $X_{\mathrm{reg}}$ admits an algebraic hyperkähler cone metric $g_{0}$. Let $Y$ be a crepant resolution of $X$. In this subsection, we show that a certain kind of openness of the twistor space in the principal twistor model holds when $X$ has an isolated singularity.
@@
Let $X$ be a conical symplectic variety with an isolated singularity, and assume that the regular locus $X_{\mathrm{reg}}$ admits an algebraic hyperkähler cone metric $g_{0}$. Let $Y$ be a crepant resolution of $X$. Let $(\mathcal{Y}(1)\to\mathcal{C}(2),\omega_{P},\tau_{P})$ be the principal twistor model given by Theorem 3.39, and let $H^{0}(\mathbb{P}^{1},\mathcal{C}(2))^{\sigma_{2}}$ denote the real vector space of real sections of the vector bundle $\mathcal{C}(2)$. For a real section $s\in H^{0}(\mathbb{P}^{1},\mathcal{C}(2))^{\sigma_{2}}$, let $Z_{s}$ be the twistor model obtained by slicing $\mathcal{Y}(1)$ along $s$. If $Z_{s}$ is the twistor space associated with a hyperkähler metric on $Y$ that is asymptotic to $g_{0}$ at infinity, then for any real section $s^{\prime}$ in a sufficiently small open neighborhood of $s$, the model $Z_{s^{\prime}}$ is also the twistor space of such a metric.
@@
    The twistor model $Z_{s^{\prime}}$ is a twistor space: Consider a relatively compact open set $U$ containing both the real section $s\in H^{0}(\mathbb{P}^{1},\mathcal{C}(2))^{\sigma_{2}}$ and the zero section. By Assertion (1), there exists a compact set $K$ containing the exceptional set $E$ on the resolution $Y$ such that the map $\Phi$ is surjective onto the set $(Y\setminus K)\times U$. On the other hand, since $Z_{s}$ is a twistor space by assumption, for the compact set $K$ on $Y$, by choosing a sufficiently small open neighborhood $U^{\prime}\subset U$ of the real section $s\in H^{0}(\mathbb{P}^{1},\mathcal{C}(2))^{\sigma_{2}}$, the map $\Phi$ is surjective onto the set $K\times U^{\prime}$. Therefore, the map $\Phi$ is surjective onto the set $Y\times U^{\prime}$. Now, the crepant resolution $Y$ is simply connected and we may assume that the neighborhood $U^{\prime}$ is also simply connected. This allows us to lift the set $Y\times U^{\prime}$ to the finite covering $\Phi$, and we can choose the lift that contains the twistor lines of $Z_{s}$. Thus, for any real section $s^{\prime}\in U^{\prime}$, the twistor model $Z_{s^{\prime}}$ is a twistor space.
@@
As a corollary of this proposition, we obtain the following result on the moduli space of hyperkähler structures with asymptotic behavior.
@@
Let $X$ be a conical symplectic variety with an isolated singularity. Assume that the regular locus $X_{\mathrm{reg}}$ admits an algebraic hyperkähler cone metric $g_{0}$. Let $Y$ be a crepant resolution of $X$. Then, the following two assertions hold:
@@
In particular, if the moduli space $\mathcal{M}$ (resp. $\mathcal{M}_{\text{met}}$) is non-empty, then we have $\dim_{\mathbb{R}}\mathcal{M}=3d$ (resp. $\dim_{\mathbb{R}}\mathcal{M}_{\text{met}}=3d-3$), where $d=\dim H^{2}(Y;\mathbb{C})$.
@@
Corollary 4.18 applies to the closures of minimal nilpotent orbits in complex semisimple Lie algebras. Let $X=\overline{\mathcal{O}_{\mathrm{min}}}$ be the closure of the minimal nilpotent orbit (cf. [3]) in a complex semisimple Lie algebra $\mathfrak{g}$. The space $X$ is a conical symplectic variety with an isolated singularity at the origin, and its regular locus $X_{\mathrm{reg}}=\mathcal{O}_{\mathrm{min}}$ admits an exact hyperkähler cone metric via Nahm’s equations, as constructed by Kronheimer [8]. As a specific example, consider the case $\mathfrak{g}=\mathfrak{sl}_{n}(\mathbb{C})$ for $n\geq 2$, where $X$ admits a crepant resolution $Y=T^{*}\mathbb{P}^{n-1}$. On this resolution, there exists the Calabi metric, which is an asymptotically conical hyperkähler metric that approaches the cone metric at infinity. In this setting, Corollary 4.18 shows that the dimension of the moduli space $\mathcal{M}_{\text{met}}$ of such asymptotic metrics is exactly $3\dim H^{2}(T^{*}\mathbb{P}^{n-1};\mathbb{C})-3=0$ (meaning the metric is unique up to scaling).
@@ 4.2.1. Metrics Constructed by Hyperkähler Quotients
    $d\langle\mu_{A},\xi\rangle(v)=\omega_{A}(v,\xi^{*}),$
    where $\langle\cdot,\cdot\rangle$ denotes the pairing between $\mathfrak{g}$ and $\mathfrak{g}^{*}$, and $\omega_{A}\coloneqq g(A\cdot,\cdot)$.
@@
By applying the theory of symplectic quotients and geometric invariant theory by Kempf-Ness and Kirwan, the following lemma is obtained (cf. [9, Theorem 3.1, Theorem 4.1]). The proof is omitted.
@@
    If $X_{\zeta}$ is non-singular, it naturally inherits a hyperkähler structure. Since the map $\pi_{I}$ preserves the induced $I$-holomorphic symplectic form, it automatically gives a crepant resolution of $X_{\zeta_{0}}$.
@@
    The hyperkähler quotient $X_{\zeta}$ is naturally diffeomorphic to a crepant resolution $Y$ of $X_{0}$: We follow the argument of Kronheimer [8]. Let $g_{\zeta}$ be the hyperkähler metric on the hyperkähler quotient $X_{\zeta}$, where we denote the point $\zeta$ by $\zeta=(\zeta_{I},\zeta_{J},\zeta_{K})\in\mathbb{R}^{3}\times\mathfrak{g}^{*}$. Since the group $G$ acts linearly on the vector space $M$, the set of points $\xi\in\mathbb{R}^{3}\times\mathfrak{g}^{*}$ such that $X_{\xi}\coloneqq\mu^{-1}(\xi)/G$ admits singularities can be written as a union of finitely many hyperplanes in the vector space $\mathbb{R}^{3}\times\mathfrak{g}^{*}$. (That is, it has a wall-chamber structure.) Therefore, by considering the $SO(3)$-action on $\mathbb{R}^{3}$ by hyperkähler rotation, we may assume that the parameter $\zeta$ is generic so that $X_{\zeta}$, $X_{\zeta^{\prime}}$, and $X_{\zeta^{\prime\prime}}$ are all non-singular, where we set $\zeta^{\prime}\coloneqq(0,\zeta_{J},\zeta_{K})$ and $\zeta^{\prime\prime}\coloneqq(0,0,\zeta_{J})$. Since these spaces are non-singular, the maps $\pi_{I},\pi_{J}$, and $\pi_{K}$ obtained in Lemma 4.24 automatically give crepant resolutions (see Remark 4.25), and we obtain a smooth map
@@
    as their composition. Note that $\pi_{I}$ and $\pi_{J}$ in the map (4.3) are diffeomorphisms. Therefore, if we let $Y\coloneqq X_{\zeta^{\prime\prime}}$ be the $K$-holomorphic crepant resolution of $X_{0}$, $X_{\zeta}$ is naturally diffeomorphic to the crepant resolution $Y$.
@@
By the above proposition, Theorem 3.39 can be applied to Nakajima quiver varieties (cf. [9]) and toric hyperkähler manifolds (cf. [1]). Note that for toric hyperkähler manifolds, this application requires the hyperkähler moment map to satisfy $\mu(0)=0$. Under this condition—which is automatically satisfied for quiver varieties by definition—one can directly verify from the explicit forms of the moment maps that the required condition $\lambda_{c}^{*}\mu=|c|^{2}\mu$ holds in both cases.
@@ 4.2.2. QALE Hyperkähler Metrics
Let $G<Sp(n)$ be a finite group, and let $X=\mathbb{C}^{2n}/G$ be the quotient space. Assume that $X$ admits a crepant resolution $Y$. We assume that any QALE hyperkähler metric on $Y$ (cf. [6, Definition 2.1]), provided it exists, is necessarily algebraic.
@@
According to Joyce [6, p.115], the above is proved by hypercomplex algebraic geometry, but the details are not described. Note that it is not known whether a QALE hyperkähler metric exists on an arbitrary crepant resolution $Y$.
@@
Note that $X$ has an isolated singularity if and only if $\dim_{\mathbb{C}}Y=2$. In this case, a QALE hyperkähler metric on $Y$ coincides with an ALE gravitational instanton. Thus, by Corollary 4.18, we deduce that the moduli space of ALE gravitational instantons has real dimension 3d33d-3, where $d=\dim H^{2}(Y;\mathbb{C})$. This result also follows immediately from the result of Kronheimer [7].